# CCaaLF: Concurrency Control as a Learnable Function

### v3

@@ Abstract.
In this paper, we propose CCaaLF (Concurrency Control as a Learnable Function), a novel learned concurrency control algorithm designed to achieve high performance across varying workloads. The algorithm is quick to optimize, making it robust against dynamic workloads. CCaaLF learns an agent function that captures a large number of design choices from existing CC algorithms. The function is implemented as an efficient in-database lookup table that maps database states to concurrency control actions. The learning process is based on a combination of Bayesian optimization and a novel graph reduction algorithm, which converges quickly to a function that achieves high transaction throughput. We compare CCaaLF against five state-of-the-art CC algorithms and show that our algorithm consistently outperforms them in terms of transaction throughput and optimization time.
@@ 1. Introduction
Concurrency control (CC) algorithms allow database transactions to exploit hardware concurrency while ensuring correctness, i.e., transaction isolation. As modern database systems are no longer disk-bounded, CC is essential for achieving high transaction processing performance. Popular CC algorithms, such as two-phase locking (2PL) and optimistic concurrency control (OCC), do not perform well across all workloads (DBLP:journals/pvldb/YuBPDS14/1000_core_cc, ; DBLP:conf/cidr/TangJE17/ACC, ). In particular, 2PL generally works better than OCC in high-contention workloads, but the reverse is true in low-contention workloads.
@@
Recent works on CC focus on combining multiple classical CC algorithms to support different workloads. The common approach follows a classify-then-assign process, in which transaction operations are classified into different types, then the most appropriate CC algorithm is selected for each type. The classification can be based on the transaction (DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/sosp/XieSLAK015/Callas, ), or based on the data (DBLP:conf/sigmod/ShangLYZC16/HSync, ; DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ). The selection of CC algorithm is either heuristic (DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) or based on learning (DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ). We note that existing works do not perform well for dynamic workloads that can drift over time. In particular, these works require manual, time-consuming changes to the classification and algorithm selection when there are workload drifts. While CormCC (DBLP:conf/usenix/TangE18/CormCC, ) and ACC (DBLP:conf/cidr/TangJE17/ACC, ) propose dynamic mechanisms that alleviate the cost of manual changes, they only support a small number of CC algorithms. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) uses an evolutionary algorithm to learn the optimal combination of CC algorithms. It learns CC actions — e.g., to wait for dependent transactions to reach specific steps, or to expose uncommitted operations — for each data access. However, it does not support general transactions, and does not work well for dynamic workloads due to the long reoptimizing time.
@@
Our goal is to design a learned CC algorithm that can achieve high performance across different workloads, and is efficient to optimize. More specifically, the algorithm can extract a high degree of concurrency from different workloads, while incurring small overhead. In addition, by quickly optimizing when the workload changes, it can perform well under dynamic workloads. Designing such a novel CC algorithm is challenging for three reasons. First, a large number of CC algorithms proposed in the literature are optimized for specific workloads. Thus, for our algorithm to perform well across different workloads, it must incorporate design choices from as many existing CC algorithms as possible. Second, to achieve high performance, the algorithm’s overhead must be much smaller than the transaction execution time. This is non-trivial because, in modern in-memory databases, transactions can be executed in microseconds (or even in nanoseconds). Third, learning the optimal action for each transaction operation is challenging because most of the learned parameters (such as the wait parameters in Polyjuice) are non-continuous. This lack of locality, as demonstrated in previous works (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), renders machine learning techniques like Policy Gradient ineffective. As a result, costly optimization strategies, such as evolutionary algorithms and random search, are used.
@@
We present CCaaLF, a novel learned CC algorithm that achieves our goal above. It addresses the first challenge by proposing a unified model of existing concurrency control algorithms. In particular, CCaaLF models a concurrency control algorithm as a learnable agent function :${\mathcal{{F}}}:\{{\mathbf{{s}}}\}\to\{{\mathbf{{a}}}\}$ that maps the database state $\mathbf{{s}}$ to a set of concurrency control actions $\mathbf{{a}}$. It captures existing classify-then-assign approaches: $\mathbf{{s}}$ represents the features used for classification, $\mathbf{{a}}$ indicates the actions of the assigned CC algorithm. In CCaaLF, a state $\mathbf{{s}}$ consists of features used by existing adaptive concurrency controls, such as data hotness (DBLP:conf/icde/LinCZ18/ASOCC, ) and number of dependent transactions in dependency graph (DBLP:journals/pvldb/TianHMS18/LDSF, ). An action $\mathbf{{a}}$ consists of a combination of conflict detection and resolution mechanisms. Instead of selecting from existing algorithms, we propose deconstructing concurrency control into a sequence of conflict detection and resolution actions, which allows for a broad range of CC designs by enabling arbitrary combinations of these actions. Our model is more general than existing adaptive concurrency control algorithms, such as Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and bLSDF (DBLP:journals/pvldb/TianHMS18/LDSF, ). The former only maps data access IDs to pipeline-related wait actions, and the latter only maps dependency set sizes to priority-related wait actions. Figure 1 illustrates how Polyjuice treats all transactions accessing the same data as equals, and consequently takes a suboptimal action that results in high abort rates. In this example, bLSDF would give higher priority to transactions with more dependent transactions. However, it does not consider transaction pipelining which enables efficient transaction execution.
Figure 1. An example of CC designs not covered in Polyjuice.
@@
CCaaLF addresses the second challenge by optimizing the cost of collecting the state $\mathbf{{s}}$ and of executing $\mathcal{{F}}$($\mathbf{{s}}$). For the former, it avoids states that are expensive to compute, e.g., hash values (DBLP:conf/icde/Su0Z21/C3, ) and conflict graphs (DBLP:conf/sigmod/WangMCYCL16/ic3, ). Instead, it uses states that can be computed in a lock-free manner, such as the number of direct dependencies and the number of executed operations. CCaaLF implements the agent function as an in-database lookup table recording the state-action pairs. This table only consumes a small amount of memory, as CCaaLF learns an efficient feature selector that reduces the input states to only hundreds of values. The whole table indeed fits in the database cache, and CC actions for an operation can be determined by a single table lookup, which only costs hundreds of CPU cycles.
@@
CCaaLF addresses the third challenge in three steps. First, it reduces the cost of evaluating the agent function. Previous works evaluate the function by loading the lookup table into the database and monitoring the system performance. This process takes seconds to minutes (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). CCaaLF adopts a Bayesian optimization approach that trains a surrogate model for predicting the system performance given different action lookup tables. By sampling from this surrogate model, it can quickly explore the search space and guide the optimization process on which agent function to evaluate next. Second, it re-formulates the original learning task of pipeline-wait actions as an optimal conflict graph learning task, which exhibits locality and therefore is amenable to common learning techniques. This is important because directly learning the pipeline-wait actions is difficult due to the lack of locality around the optimal combination of them. In particular, a small modification to an action could lead to dependency cycles and significant performance reduction. CCaaLF learns the table indirectly by solving a conflict graph learning problem, for which we design an efficient algorithm similar to genetic search. Third, CCaaLF speeds up the learning process by prioritizing the actions that yield higher improvements. Specifically, we prioritize fine-tuning wait-related actions that can bring higher performance improvements over actions like timeouts and back-off durations.
@@
  • We present CCaaLF, a novel learned concurrency control algorithm that achieves high performance across diverse workloads. The algorithm is also quick to optimize, which enables it to adapt quickly to workload drifts.
@@
  • We conduct extensive evaluation of CCaaLF, comparing it against five state-of-the-art CC algorithms. Experimental results demonstrate that CCaaLF consistently outperforms the baselines both in throughput and in optimization time.
The remainder of the paper is structured as follows. Section 2 discusses the design space of concurrency control algorithms. Section 3 provides an overview of CCaaLF. Section 4 details how CCaaLF performs fine-grained concurrency control via an optimized action lookup table. Section 5 describes the learning process. Section 6 presents the experimental evaluation. Section 7 discusses related works, and Section 8 concludes.
@@ 2. Concurrency Control Design Space
In this section, we first describe the design space of concurrency control (CC) algorithms. We discuss how existing algorithms fit into the space, and identify the gaps that represent new opportunities.
A transaction is modeled as a sequence of data access operations, each being either a read or a write. CC algorithms allow transactions to be executed concurrently and correctly (DBLP:conf/sigmod/AdyaGLM95/occ, ; DBLP:conf/sosp/TuZKLM13/silo, ; DBLP:reference/db/Fekete18c/ssi, ; DBLP:conf/sigmod/LimKA17/cicada, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:journals/pvldb/WangK16/mocc, ; DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:conf/eurosys/MuAS19/DRP, ; DBLP:journals/pvldb/FaleiroAH17/EWV, ; 2PL_no_wait, ; DBLP:journals/tods/RosenkrantzSL78/2PL_wait_die, ; DBLP:journals/pvldb/BarthelsMTAH19/2PL_bound_wait, ; DBLP:journals/tods/HsuZ92/cautious_wait, ; DBLP:conf/icde/FranaszekRT91/wait_depth_limit, ; DBLP:conf/sigmod/ShengTZP19/abort_prediction, ; DBLP:conf/pods/CareyKL90/half_half, ; DBLP:conf/vldb/MoenkebergW92/opt_DMP, ; DBLP:conf/icde/MoenkebergW91/conflict_ratio, ; DBLP:conf/sigmod/ThomsonDWRSA12/Calvin, ; DBLP:journals/pvldb/KallmanKNPRZJMSZHA08/HStore, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sosp/XieSLAK015/Callas, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:journals/pvldb/LuYCM20/Aria, ; DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/sigmod/ShangLYZC16/HSync, ; DBLP:conf/icde/Su0Z21/C3, ; NeurDB, ). To ensure correctness, a CC algorithm employs CC actions — algorithmic steps that ensure safe interleaving of data access operations among concurrent transactions. In the following, we describe a design space of CC actions consisting of two design dimensions: conflict detection and conflict resolution. A CC algorithm can use one or multiple actions drawn from this space.
@@
2.1. Conflicts Detection
@@
We consider two transactions $T_{1}$ and $T_{2}$ executing concurrently. They both perform read and write on data objects A and B. Figure 2 illustrates examples where $T_{2}$ has Read-Write (RW) conflicts with $T_{1}$, and $T_{2}$ uses different conflict detection approaches.
@@
Detect critical. The CC algorithm avoids the cost of aborts by identifying conflicts that are likely to cause aborts based on heuristic rules. They are called critical conflicts, which may lead to validation failure or to dependency cycles. Upon detection, the algorithm immediately aborts the transaction, e.g., ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ) and SSI (DBLP:reference/db/Fekete18c/ssi, ), or reorders operations to allow conflicting transactions to commit together, e.g., Bamboo (DBLP:conf/sigmod/GuoWYY21/bamboo, ) and IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ). In Figure 2(b), $T_{2}$ detects that $T_{1}$ updated the previously read A during early validation and immediately aborts without executing to the end. In Figure 2(c), $T_{2}$ detects that the immediate read of B would form a dependency cycle between $T_{1}$ and $T_{2}$. It then reorders its read of B to occur after $T_{1}$’s write of B, thus breaking the cycle.
@@
Detect all. The CC algorithm detects all conflicts that may lead to transaction aborts. The classical 2PL and its variants (2PL_no_wait, ; DBLP:journals/tods/RosenkrantzSL78/2PL_wait_die, ; DBLP:journals/pvldb/BarthelsMTAH19/2PL_bound_wait, ) adopt this approach. Figure 2(d) shows an example in which this approach causes $T_{2}$ to block for a long time, which increases transaction latency. Furthermore, the detection algorithm introduces some overhead because it has to find all conflicts.
2.2. Conflicts Resolution
@@
Once a conflict is detected, the concurrency control algorithm resolves it by coordinating accesses from the conflicting transactions. We identify two design choices for conflict resolution: timeout, and priority.
@@
Timeout. When the CC algorithm detects a conflict, it blocks the current transaction for a specific duration (a timeout), waiting for the conflicting operations to end in this duration, and aborts it when the timeout is exceeded. A timeout of 00 means that the transaction aborts immediately (2PL_no_wait, ), which assumes pessimistically that the transaction encountering the conflict will eventually abort. A large timeout, or timeout of $\infty$, means the transaction is blocked. In particular, the transaction is either put on a wait list associated with the data object, or is entering a spinning loop until the data object is available. This approach is implemented in lock-based and pipeline-based CC algorithms, e.g., 2PL and IC3. The blocking overhead can be significant, especially when the transaction eventually aborts.
@@
Priority. When multiple transactions are blocked on the same data object, the algorithm may order them based on their priorities. When the object becomes available, the highest-priority transaction is woken up first. Most existing concurrency control algorithms consider no priority (DBLP:conf/sosp/TuZKLM13/silo, ; DBLP:journals/tods/HsuZ92/cautious_wait, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ), or simple FIFO priority (DBLP:journals/pvldb/BarthelsMTAH19/2PL-BW, ). Recent works heuristically assign priorities to conflicting transactions (DBLP:conf/eurosys/HuangMW17/VProfiler, ; DBLP:conf/sigmod/HuangMSW17/VATS, ; DBLP:journals/pvldb/TianHMS18/LDSF, ). For example, bLDSF (DBLP:journals/pvldb/TianHMS18/LDSF, ) gives higher priorities to transactions with higher numbers of dependent transactions. Other algorithms (DBLP:conf/eurosys/HuangMW17/VProfiler, ; DBLP:conf/sigmod/HuangMSW17/VATS, ) consider transaction age and the number of held locks when assigning priorities.
@@
Table 1. Comparison of state-of-the-art concurrency control algorithms and CCaaLF in the design space. Timeout is omitted because it is supported by all algorithms.
@@
OCC (DBLP:conf/sigmod/AdyaGLM95/occ, ), Silo (DBLP:conf/sosp/TuZKLM13/silo, )
2PL (2PL_no_wait, ), TSO (DBLP:books/aw/BernsteinHG87/tso1, ; DBLP:conf/icde/LeisK014/tso2, ), MVCC (wu2017empirical/mvcc, )
EWV (DBLP:journals/pvldb/FaleiroAH17/EWV, ), IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, )
bLDSF (DBLP:journals/pvldb/TianHMS18/LDSF, ), VATS (DBLP:conf/sigmod/HuangMSW17/VATS, )
MCC (DBLP:conf/sosp/XieSLAK015/Callas, ), Tebaldi (DBLP:conf/sigmod/SuCDAX17/Tebaldi, )
ACC (DBLP:conf/cidr/TangJE17/ACC, ), C3 (DBLP:conf/icde/Su0Z21/C3, ), CormCC (DBLP:conf/usenix/TangE18/CormCC, )
ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ), Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, )
CCaaLF
Figure 3. CCaaLF’s concurrency control (CC) execution workflow and function optimization workflow.

### v4

# Modeling Concurrency Control as a Learnable Function

@@ Abstract.
In this paper, we propose NeurCC, a novel learned concurrency control algorithm that achieves high performance across diverse workloads. The algorithm is quick to optimize, making it robust against dynamic workloads. It learns a function that captures a large number of design choices from existing CC algorithms. The function is implemented as an efficient in-database lookup table that maps database states to concurrency control actions. The learning process is based on a combination of Bayesian optimization and a novel graph reduction search algorithm, which converges quickly to a function that achieves high transaction throughput. We compare NeurCC against five state-of-the-art CC algorithms and show that it consistently outperforms the baselines both in transaction throughput and in optimization time.
conference: ; ; copyright: none
@@ 1. Introduction
Concurrency control (CC) algorithms allow database transactions to exploit hardware concurrency while ensuring correctness, i.e., transaction isolation. As modern database systems are no longer disk-bound, CC is essential for achieving high transaction processing performance. Classical CC algorithms, such as two-phase locking (2PL) and optimistic concurrency control (OCC), do not perform well across all workloads (Yu et al., 2014; Tang et al., 2017). In particular, 2PL typically outperforms OCC for high-contention workloads, but the reverse is true for low-contention workloads.
@@
Recent works on CC combine multiple classical CC algorithms to support different workloads. They adopt the classify-then-assign approach, in which transaction operations are classified into different types, then the most appropriate CC algorithm is selected for each type. The classification can be based on the transaction (Tai and Meyer, 1996; Su et al., 2021, 2017; Xie et al., 2015), or on the data (Shang et al., 2016; Tang et al., 2017; Lin et al., 2018; Tang and Elmore, 2018; Wang et al., 2021). The selection of CC algorithm is either heuristic (Su et al., 2021; Tai and Meyer, 1996; Lin et al., 2018; Tang and Elmore, 2018; Su et al., 2017; Hong et al., 2025) or based on learning (Tang et al., 2017; Wang et al., 2021). We note that these works do not perform well for dynamic workloads that can drift over time. In particular, they require manual, time-consuming changes to the classification and algorithm selection when there are drifts. CormCC (Tang and Elmore, 2018) and ACC (Tang et al., 2017) propose dynamic mechanisms that reduce the cost of manual changes, but they only support a small number of CC algorithms. Polyjuice (Wang et al., 2021) uses an evolutionary algorithm to learn the optimal combination of CC algorithms. It learns CC actions — e.g., to wait for dependent transactions to reach specific steps, or to expose uncommitted operations — for each data access. However, it does not support general transactions and does not work well for dynamic workloads due to the long optimization time.
@@
Our goal is to design a learned CC algorithm that can achieve high performance across diverse workloads, and is efficient to optimize. More specifically, the algorithm can extract a high degree of concurrency from different workloads, while incurring small overhead. The short optimization time allows it to perform well under dynamic workloads. There are three challenges in designing this algorithm. First, a large number of CC algorithms proposed in the literature are optimized for specific workloads. Thus, for our algorithm to perform well across different workloads, it must incorporate design choices from as many existing CC algorithms as possible. Second, to achieve high performance, the algorithm’s overhead must be smaller than the transaction execution time. This is non-trivial because, in modern in-memory databases, transactions can be executed in microseconds (or even in nanoseconds). Third, learning the optimal action for each transaction operation is challenging because most of the learned parameters (such as the wait parameters in Polyjuice) are non-continuous. This lack of locality renders machine learning techniques such as Policy Gradient (Sutton et al., 1999) ineffective, thus algorithms like Polyjuice resort to costly optimization strategies such as evolutionary algorithms and random search.
@@
We present NeurCC, a novel learned CC algorithm that achieves the goal above. It addresses the first challenge by proposing a unified model of existing concurrency control algorithms. In particular, NeurCC models a concurrency control algorithm as a learnable function ${\mathcal{{F}}}:\{{\mathbf{{s}}}\}\to\{{\mathbf{{a}}}\}$ that maps the database state $\mathbf{{s}}$ to a set of concurrency control actions $\mathbf{{a}}$. It captures the existing classify-then-assign approaches: $\mathbf{{s}}$ represents the features used for classification, $\mathbf{{a}}$ indicates the actions of the assigned CC algorithm. In NeurCC, a state $\mathbf{{s}}$ consists of features used by existing adaptive concurrency controls, such as data hotness (Lin et al., 2018) and the number of dependent transactions in the dependency graph (Tian et al., 2018). An action $\mathbf{{a}}$ consists of a combination of conflict detection and resolution mechanisms. Instead of selecting from existing algorithms, NeurCC decomposes concurrency control into a sequence of conflict detection and resolution actions, which allows for a wide range of CC designs by enabling arbitrary combinations of these actions. The model is more general than existing adaptive concurrency control algorithms, such as Polyjuice (Wang et al., 2021) and bLDSF (Tian et al., 2018). The former only maps data access IDs to pipeline-wait actions, and the latter only maps dependency set sizes to priority-related wait actions. Figure 1 illustrates the CC designs covered by Polyjuice and bLDSF. Specifically, Polyjuice treats all transactions accessing the same data as equals, which could result in suboptimal actions with high abort rates. bLDSF gives higher priority to transactions with more dependent transactions, but it does not consider transaction pipelining, which is essential for efficient transaction execution.
Figure 1. Example of CC designs covered by different algorithms.
Diagram showing concurrency control designs not covered by the Polyjuice system.
@@
NeurCC addresses the second challenge by making the collection of the state $\mathbf{{s}}$ and the execution of $\mathcal{{F}}$($\mathbf{{s}}$) highly efficient. Specifically, it excludes states that are expensive to compute, e.g., hash values (Su et al., 2021) and conflict graphs (Wang et al., 2016), and instead uses states that can be computed in a lock-free manner, such as the number of direct dependencies and the number of executed operations. NeurCC implements $\mathcal{{F}}$ as an in-database lookup table for the state-action pairs. This table only consumes a small amount of memory, as NeurCC learns an efficient feature selector that reduces the input states to only hundreds of values. The whole table fits in the database cache, and the CC actions for an operation can be determined by a single table lookup, which only costs hundreds of CPU cycles.
@@
NeurCC addresses the third challenge in three steps. First, it reduces the cost of evaluating the function’s performance. Previous works evaluate the function’s performance by loading the lookup table into the database and monitoring the system performance. This process takes seconds to minutes (Wang et al., 2021). NeurCC adopts a Bayesian optimization approach that trains a surrogate model for predicting the system performance given an action lookup table. By sampling from this surrogate model, it can quickly explore the search space and guide the optimization process to select the next function to evaluate. Second, NeurCC re-formulates the original learning task of pipeline-wait actions as an optimal conflict graph learning task that exhibits locality and therefore is amenable to common learning techniques. This reformulation is important since learning the pipeline-wait actions directly is challenging due to the lack of locality around the optimal combination of them. In particular, a small modification to an action could lead to dependency cycles and significant performance reduction. NeurCC uses an efficient algorithm similar to genetic search for this conflict graph learning task. Third, NeurCC speeds up the learning process by prioritizing the actions that yield higher improvements. Specifically, it prioritizes fine-tuning wait-related actions that can bring higher performance improvements over actions like timeouts and back-off durations.
@@
  • We present NeurCC, a novel learned concurrency control algorithm that achieves high performance across diverse workloads. The algorithm is quick to optimize, allowing it to adapt quickly to workload drifts.
@@
[…]
The remainder of the paper is structured as follows. Section 2 discusses the design space of concurrency control algorithms. Section 3 provides an overview of NeurCC. Section 4 details how NeurCC performs fine-grained concurrency control via an optimized action lookup table. Section 5 describes the learning process. Section 6 outlines the deployment workflow. Section 7 presents the experimental evaluation. Section 8 discusses related works, and Section 9 concludes.
@@ 2. Concurrency Control Design Space
In this section, we first describe the design space of concurrency control (CC) algorithms. We then discuss how existing algorithms fit into the space, and identify the gaps representing new opportunities.
A transaction is modeled as a sequence of data access operations, each being either a read or a write. CC algorithms allow transactions to be executed concurrently and correctly (Adya et al., 1995; Tu et al., 2013; Fekete, 2018; Lim et al., 2017; Wang et al., 2016; Wang and Kimura, 2016; Guo et al., 2021; Mu et al., 2019; Faleiro et al., 2017; Chesnais et al., 1983; Rosenkrantz et al., 1978; Barthels et al., 2019; Hsu and Zhang, 1992; Franaszek et al., 1991; Sheng et al., 2019; Carey et al., 1990; Mönkeberg and Weikum, 1992, 1991; Thomson et al., 2012; Kallman et al., 2008; Su et al., 2017; Tang and Elmore, 2018; Xie et al., 2015; Wang et al., 2021; Lu et al., 2020; Tang et al., 2017; Lin et al., 2018; Shang et al., 2016; Su et al., 2021; Zhao et al., 2024). To ensure correctness, a CC algorithm employs CC actions — algorithmic steps that ensure safe interleaving of data access operations among concurrent transactions. In the following, we describe the design space of CC actions consisting of two design dimensions: conflict detection and conflict resolution. A CC algorithm can use one or multiple actions from this space.
@@
Diagram showing different conflict detection approaches.
2.1. Conflict Detection
@@
We consider two transactions $T_{1}$ and $T_{2}$ executing concurrently. They both perform read/write on data object A and B. Figure 2 illustrates examples where $T_{2}$ has Read-Write (RW) conflicts with $T_{1}$, and $T_{2}$ uses different conflict detection approaches.
@@
Detect critical. The CC algorithm avoids the cost of aborts by identifying conflicts that are likely to cause aborts based on heuristic rules. They are called critical conflicts, which may lead to validation failure or to dependency cycles. Upon detection, the algorithm immediately aborts the transaction (early validation), e.g., ASOCC (Lin et al., 2018) and SSI (Fekete, 2018), or reorders operations to allow conflicting transactions to commit together (pipeline-wait), e.g., Bamboo (Guo et al., 2021) and IC3 (Wang et al., 2016). In Figure 2(b), $T_{2}$ detects that $T_{1}$ updated the previously read A during early validation and immediately aborts without executing to the end. In Figure 2(c), $T_{2}$ detects that the immediate read of A would form a dependency cycle between $T_{1}$ and $T_{2}$. It then reorders its read of A to occur after $T_{1}$’s write of A, thus breaking the cycle.
@@
Detect all. The CC algorithm detects all conflicts that may lead to transaction aborts. The classical 2PL and its variants (Chesnais et al., 1983; Rosenkrantz et al., 1978; Barthels et al., 2019) adopt this approach. Figure 2(d) shows an example in which this approach causes $T_{2}$ to block for a long time, which increases transaction latency. Furthermore, the detection algorithm introduces some overhead because it has to find all conflicts.
2.2. Conflict Resolution
@@
Once a conflict is detected, the concurrency control algorithm resolves it by coordinating accesses from the conflicting transactions. We identify two design choices for conflict resolution: timeout and priority.
@@
Timeout. When the CC algorithm detects a conflict, it blocks the current transaction for a specific duration (a timeout), waiting for the conflicting operations to end in this duration, and aborts it when the timeout is exceeded. A timeout of 0 means that the transaction aborts immediately (Chesnais et al., 1983), which assumes pessimistically that the transaction will eventually abort. A large timeout, or timeout of $\infty$, means the transaction is blocked. In particular, the transaction is either put on a wait list associated with the data object, or is entering a spinning loop until the data object is available. This approach is implemented in lock-based and pipeline-wait-based CC algorithms, e.g., 2PL and IC3. The blocking overhead can be significant, especially when the transaction eventually aborts.
@@
Priority. When multiple transactions are blocked on the same data object, the algorithm may order them based on their priorities. When the object becomes available, the highest-priority transaction is woken up first. Most existing CC algorithms either do not consider priority (Tu et al., 2013; Hsu and Zhang, 1992; Wang et al., 2021), or use simple FIFO or timestamp priority (Barthels et al., 2019; Rosenkrantz et al., 1978; Guo et al., 2021; Zhang et al., 2025). Recent works make heuristic priority assignment to conflicting transactions (Huang et al., 2017b, a; Tian et al., 2018; Ye et al., 2023). For example, bLDSF (Tian et al., 2018) gives higher priorities to transactions with higher numbers of dependent transactions. Other algorithms (Huang et al., 2017b, a) consider transaction age and the number of held locks when assigning priorities.
@@
Table 1. Comparison of state-of-the-art concurrency control algorithms and NeurCC in the design space. Timeout is omitted because it is supported by all algorithms. Tuned priority refers to either heuristic or learned priority assignments.
@@
OCC (Adya et al., 1995), Silo (Tu et al., 2013)
2PL (Chesnais et al., 1983), TSO (Bernstein et al., 1987; Leis et al., 2014), MVCC (Wu et al., 2017)
Bamboo (Guo et al., 2021), Rebirth-Retire (Zhang et al., 2025), EWV (Faleiro et al., 2017), IC3 (Wang et al., 2016)
bLDSF (Tian et al., 2018), VATS (Huang et al., 2017a)
Polaris (Ye et al., 2023)
MCC (Xie et al., 2015), Tebaldi (Su et al., 2017)
ACC (Tang et al., 2017), TxnSails (Zhuang et al., 2025), C3 (Su et al., 2021), CormCC (Tang and Elmore, 2018), HDCC (Hong et al., 2025)
ASOCC (Lin et al., 2018), Polyjuice (Wang et al., 2021)
NeurCC